Suppose that an airline accepted 12 reservations for a commuter plane with 10 seats.
They know that 7 reservations went to regular commuters who will show up for sure.
The other 5 passengers will show up with a 50% chance, independently of each other.
Find the probability that the
flight will be overbooked.
1
Expert's answer
2020-06-08T18:39:46-0400


Let, X = the random variable denoting the number of non-regular passengers who will show up, X = 0, 1, 2, 3, 4, 5.


Since, the chance of each of these passengers will show up is 50% i.e. 0.5 and is independent of each other


Therefore, X ~ bin(n = 5, p = 0.5)


The p.m.f. of X is given by,


P(X = x) = "\begin{cases}
 \dbinom{5}{x}(0.5)^x(1-0.5)^{5-x} &\text{for } x=0,1,...,5 \\
 0 &\text{otherwise }
\end{cases}"


The commuter plane has 10 seats and out of 12 reservations, airline knows that 7 reservations went to regular commuters who will show up for sure. So, there are 3 seats remaining and 5 reservations. If more than 3 passengers out of 5 reservations show up, then the flight will be overbooked.


"\therefore" The probability that the flight will be overbooked


P(X > 3)


= P(X = 4) + P(X = 5)


= "\dbinom{5}{4}(0.5)^4(1-0.5)^{5-4}+\dbinom{5}{5}(0.5)^5(1-0.5)^{5-5}"


= 0.15625 + 0.03125


= 0.1875


Answer: The probability that the flight will be overbooked is 0.1875.